ind × Timer Notes Inductance of Solenoid sure you are radius and diameter correctly, etc. O Type here to
media%2F579%2F579ac733-9090-4d6f-bff0-3b
Answer #1

First we will find number of turns,

N = n*L

where n = 1 / 7.8e-3

n = 128

N = 128*2.6 = 333 turns

Now, L (Inductance) = uo*N2*A / Length

L = 4*PI*10-7*333*333*7.0685E-4 / 2.6

L = 3.788e-5 H

L/m = 3.788e-5 / 2.6

L/m = 1.45e-5 H/m